my mother works as a tutor and a babysitter. she earns 200.00 per hour as a tutor and 100.00 per hour as a babysitter. she renders 12 hours a week. she works to get at least 2,000.00. write the system of inequalities that will represent this situation. find out at least the possible number of hours my mother works as a tutor and a babysitter in a week, by graphing


1
Expert's answer
2021-12-03T07:49:13-0500

Suppose that mother works "x" hours as a tutor and "y" hours as a babysitter in a week. Then


"x+y=12"

"200x+100y\geq2000"

"x\geq 0, y\geq 0"

Then

"\begin{matrix}
 x+y=12 \\
 2x+y\geq20\\
x\geq 0, y\geq 0
\end{matrix}"


Mother must work at least "8" hours to get at least 2,000.00.